PERFORMANCE OF A CONTINUOUS SUPERSONIC EXPANSION DISCHARGE NOZZLE EVALUATED BY LASER-INDUCED FLUORESCENCE SPECTROSCOPY

\underline{KYLE N. CRABTREE}, CARRIE A. KAUFFMAN, Department of Chemistry, University of Illinois, Urbana, IL, 61801; BENJAMIN J. MCCALL, Departments of Chemistry and Astronomy, University of Illinois, Urbana, IL, 61801.

We have recently constructed a prototype continuous supersonic expansion discharge nozzle for the production of rotationally cold molecular ions. To assess the performance of this source, we have employed laser-induced fluorescence spectroscopy to measure the rotational temperature distributions of I_{2} and N_{2}^{+} as a function of position within the expansion. These measurements are performed on the B^{3}Π_{0_{u}^{+}} − X^{1}Σ_{g}^{+} (6-0), (8-1), (10-2), and (12-3) bands of I_{2} at 608 nm and the A^{2}Π_{u} − X^{2}Σ_{g}^{+} (4-0) band of N_{2}^{+} at 614 nm using a tunable cw dye laser. The temperature distributions obtained act as a feedback mechanism that aids in refinement of the source design with the aim of optimizing the densities and temperatures of the species within the expansion.